1. Work: u x2 +12=2y 24-12 12. 4x2-26-218x) 4x +4 168 1718 4x7-16xilo Vibe lac the FIGL 4 (0)18) 11 4.2.16 49-4 16 (x,y)=
Answer #1

Ans: 42²+12&ay. 4a². ay 19. un = a(4-6) 2² = 8 (4-6) X₂ ma (4-6) about the Symmetry about y-anis asu. ... o 0 = á (7-67 a 2.0